OpenStudy (anonymous):

Write an equation of the line containing the given point and parallel to the given line. Express as y=mx+b (-4,5); 5x=4y+3

OpenStudy (anonymous):

so ur equation in slope intercept form is gonna be y = (5/4)x - (3/4) parallel means the same slope... so in ur new equation parallel, you have y = (5/4)x + b they give u a coordinate - plug and chug 5 = (5/4)(-4) + b 10 = b so ur answer is y = (5/4)x + 10

OpenStudy (anonymous):

Thank you!